# Actor-Critic Model Predictive Control

### v5

@@ IV-C Ablation study: robustness to disturbances
(a) Simulation results
(b) Real-world results
@@
IV-D Perception Aware Flight
Additionally, we train our approach in the task of flying in a circle while keeping a certain point in the center of the camera frame, similar to the approach in [46]. Given an interest point in the world frame (which is marked as an orange star in Fig. 5), we minimize the angle between the Z-axis of the camera and the line that joins the center point of the camera with the objective. This reward term is then summed to the previously presented progress reward term, which incentivizes the drone to move through the gates. In Fig. 5, we show how both AC-MLP and AC-MPC approaches can learn this behavior. The black arrows represent the direction of the camera Z-axis. Since yaw control effectiveness is the lowest for a quadrotor – a large amount of actuation is needed for a small change in yaw – this task poses a competing reward problem, where if the drone moves faster, it will necessarily be at the expense of losing perception awareness. This is the reason behind the unnatural shapes that emerge, shown in Fig. 5.
IV-E Real-world Deployment
We test our approach in the real world with a high-performance racing drone. We deploy the policy with two different race tracks: Circle track and SplitS track. We use the Agilicious control stack [45] for the deployment. The main physical parameters and components of this platform are referred to [14], under the name 4s drone. Fig. 4 illustrates the trajectories that flew in both simulation and the real world: our policy transfers to the real world without fine-tuning.
Figure 5: The proposed Actor-Critic Model Predictive Controller trained for the task of agile perception aware flight. The star in the middle of the track is the point to be kept in the middle of the camera frame.
The real-world experiments are also shown in the supplementary video.
IV-F Training time and inference time
In Table II we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths.
TABLE II: Solve times and inference times for different variations.
@@ IV-F Training time and inference time
V Discussion and Conclusion
This work presented a new learning-based control framework that combines the advantage of differentiable model predictive control with actor-critic training. We showed that our method can tackle challenging control tasks with a highly nonlinear and high-dimensional quadrotor system, and achieves robust control performance for agile flight. Additionally, our approach achieved zero-shot sim-to-real transfer, demonstrated by successfully controlling a quadrotor at velocities of up to 14 m/s in the physical world.
However, there are some limitations to be mentioned and to be improved in the future. First, an analytic model of the system is required for the differentiable MPC block, which limits our approach to mainly systems where the dynamics are known beforehand. Furthermore, training AC-MPC takes significantly longer than AC-MLP (see Table II), due to the fact that an optimization problem needs to be solved for both the forward and the backward pass through the actor network. In fact, there are open-source libraries [37, 38] that are recently evolving and implementing more efficient versions of differentiable MPC. Another limitation is that the differentiable MPC controller does not support state constraints. This could be addressed by adding state constraints to the implementation. Showcasing the approach in different tasks with different robots is another future direction.
We believe that the proposed method represents an important step in the direction of generalizability and robustness in RL. It demonstrates that modular solutions that combine the best of learning-centric and model-based approaches are becoming increasingly promising. Our approach potentially paves the way for the development of more robust RL-based systems, contributing positively towards the broader goal of advancing AI for real-world robotics applications.
@@ Acknowledgments
We would like to thank Brandon Amos, for sharing his insights regarding the differentiable MPC code, and Jiaxu Xing for the insightful discussions.

### v6

@@ V-B Is our approach robust to external disturbances?
V-C Is our approach robust to changes in the dynamics?
Since our approach integrates a model-based controller that explicitly accounts for system dynamics, the question arises of whether AC-MPC is sensitive to dynamic parameters changes when compared to AC-MLP. To investigate this, we conduct a series of experiments in which key dynamic parameters are systematically varied. Both policies are trained on the SplitS track using nominal parameter values, after which they are evaluated across a range of perturbed parameters for both AC-MLP and AC-MPC. To quantify the robustness of these policies, we deploy them from 64 different positions, equally spread on a cube of $0.5\,m$ of side length. Success is defined as the percentage of trajectories that successfully navigate through all gates of the SplitS track. Fig. 6 presents the success rate as a function of variations in mass (up to $+27$%), in XY inertia (up to ±$\pm 80$%) , and body rate limits (up to ±$\pm 40\%$).
The results demonstrate that the AC-MPC exhibits superior generalization to dynamic parameter variations compared to the AC-MLP. This improved robustness is attributed to the incorporation of a dynamic model within the differentiable MPC block, which allows for adjustments in response to perturbed dynamics during deployment without retraining.
V-D What do we gain from Model-Predictive Value Expansion?
In this section, we conduct empirical experiments to evaluate the effect of the AC-MPC extended with Model-Predictive Value Expansion algorithm (introduced in Section III-H) in terms of training performance. To isolate the effect of MPVE and enable clear reward evaluation, we employ a straightforward hovering task for the drone. Here, the drone starts at a random position and orientation within a 1x1x1 meter cube around a designated equilibrium point (0, 0, 5 meters). The objective is to stabilize the drone at this point with zero velocity and a perfectly upright orientation (hover). We compare the training performance of AC-MPC-MPVE against AC-MLP. Both algorithms are trained with a fixed sample budget of 50 steps, corresponding to roughly 2 million samples collected from the environment. Because the MPVE algorithm primarily affects the learning of the critic, which uses the Generalized Advantage Estimation (GAE) algorithm [127], we explore the impact of our extension for various values of the temporal difference parameter $\lambda$ in TD($\lambda$). The parameter $\lambda$ determines the trade-off between bias and variance when estimating returns. A lower value of $\lambda$ favors lower variance but higher bias by relying more on immediate rewards, while higher values reduce bias but increase variance by incorporating more long-term information. To investigate this, we experiment with six different values for $\lambda$: 0.0, 0.2, 0.4, 0.6, 0.8, and 1.0. For each value, we repeat the training process three times with different random seeds.
Our results (visualized in Fig. 7) demonstrate that the sample efficiency gains achieved by MPVE are more significant for lower lambda values. This can be explained by the way the value function leverages data from the rollout buffer. Higher lambda values incorporate data from a larger portion of the rollout, increasing the influence of Monte Carlo estimates and reducing the relative impact of the information provided by MPVE predictions. In contrast, lower lambda values lead to stronger benefits from MPVE. This suggests that our algorithm is particularly advantageous for scenarios where the value function is trained with TD(0). This approach is often preferred when memory limitations exist (as it avoids storing the entire rollout buffer) or for non-episodic tasks.
In Fig. 8, we showcase the evolution of the episode returns for AC-MLP and AC-MPC-MPVE under two different values of $\lambda$ and for the same task of drone stabilization. In both settings, our approach substantially outperforms the baseline in terms of asymptotic performance and sample efficiency, highlighting the benefits of model-predictive value expansion.
Figure 7: Reward at 50 steps of training for AC-MLP and AC-MPC(MPVE), depending on the $\lambda$ used in the TD($\lambda$). The policies are trained with a sample budget of 50 steps and the final reward is recorded, for different values of $\lambda$, for both AC-MPC(MPVE) and AC-MLP.
Figure 8: Reward evolution for the cases of $\lambda=0.0$ and $\lambda=0.6$, for AC-MLP and for AC-MPV(MPVE).
V-E Can we interpret the internal workings of AC-MPC?
This section investigates a potential connection between the critic network’s output, $V(\bm{x})$, and the learned cost matrix $Q(\bm{x})$. To simplify the analysis, we introduce three key assumptions. First, we assume that the observations $s_{k}$ are equivalent to the state of the system $x_{k}$. Second, the reward function is assumed to be quadratic in the state, expressed as $r(\bm{x})=-\bm{x}^{T}\bm{Q}\bm{x}$. This formulation captures the objective of stabilizing the system around a hover point. And third, we restrict the Model Predictive Control (MPC) horizon length to $T=2$. Under these simplifications, if we had a linear system with no constraints, the problem becomes equivalent to a Linear Quadratic Regulator (LQR) problem. In LQR, the value function is the solution to the Riccati equation, which is inherently quadratic with respect to the state.
Given the focus on drone dynamics and a stabilization task, we expect the system to primarily operate near hover states. This implies that the learned value function should also be approximately quadratic. The central question we aim to address here is whether a relationship exists between the Hessian (the matrix containing all second-order partial derivatives) of this value function and the learned cost matrix, $Q$. The general infinite horizon problem for an MPC can be written as:
$\pi^{*}(x_{0})$ $=\operatorname*{argmin}_{u}\sum_{k=0}^{\infty}l(x_{k},u_{k})$
$=\operatorname*{argmin}_{u}\sum_{k=0}^{N-1}l(x_{k},u_{k})+\sum_{k=N}^{\infty}l(x_{k},u_{k})$
$=\operatorname*{argmin}_{u}\sum_{k=0}^{N-1}l(x_{k},u_{k})-V^{*}(x_{N})$
$=\operatorname*{argmin}_{u}\sum_{k=0}^{N-1}l(\tau_{k})-V^{*}(f(\tau_{N-1})),$
where $\tau_{k}=\begin{bmatrix}x_{k}$, $u_{k}\end{bmatrix}$ is a state-action pair at time $k$ and the value function term is negative because we have defined it in the context of RL, where it is maximized.
Figure 9: Empirical probing of the correlation between the Hessian and the gradient of the learned critic $V(x)$ and the learned differentiable MPC cost terms $Q_{N}$ and $p_{N}$. (A): We train a linear classifier to infer $Q_{N}$ and $p_{N}$ from the Hessian of $V(x)$ and show the prediction from different 2D projections of $Q_{N}$ and $p_{N}$. We show how the groundtruth can be predicted with a high accuracy by only using a small linear classifier, which indicates that the data learned by the critic and $Q_{N}$ and $p_{N}$ are highly correlated. (B): Top: 3D plot of $\tau_{N}^{T}Q_{N}\tau_{N}+\tau_{N}^{T}p_{N}$ for the state dimensions of $v_{x}$ and $v_{y}$ after training. Bottom: 3D plot of $V(x)$ for the same dimensions. Apart from the change in sign and in scale, the similarity between both shapes indicate a high correlation between the value function and the learned cost parameters.
On the other hand, the quadratized problem that we solve in the differentiable MPC block has the following form:
$\sum_{k=0}^{N}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}^{T}Q_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}+p_{k}\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix}.$ (11)
This motivates the investigation of a potential relationship between the value function $V(x)$ learned by the critic and the quadratic ($Q_{N}$) and linear ($p_{N}$) terms learned by the actor within the differentiable MPC framework. In particular, we aim to identify a connection between the first and second derivatives of the value function and $p_{N}$ and $Q_{N}$, respectively. To achieve this, we can introduce the second-order Taylor Series approximation of the value function around the state-action pair $\tau_{N-1}$:
$V(f(\tau_{N-1}))\bigg{\rvert}_{\tau_{N-1}=\bm{d}}$ $\approx V(f(\bm{d}))$
$\quad+\tau_{N-1}^{T}\frac{\partial^{2}V(f(\tau_{N-1}=\bm{d}))}{\partial\tau_{N-1}^{2}}\tau_{N-1}$
$\quad+\frac{\partial V(f(\tau_{N-1}=\bm{d}))}{\partial\tau_{N-1}}\tau_{N-1}$
$=V(f(\bm{d}))+\tau_{N-1}^{T}H_{V}\tau_{N-1}+\Delta_{V}\tau_{N-1},$
where $f(\cdot)$ represents the system dynamics function, $\bm{d}$ denotes a specific state-action pair, $H_{V}$ is the Hessian of the value function evaluated at $\bm{d}$, $\Delta_{V}$ represents the first-order term of the Taylor expansion. Because we need to compute the gradient and the Hessian of the value function composed with the dynamics, here are the expressions for $\Delta_{V}$ and $H_{V}$.
$\Delta_{V}$ $=\frac{\partial}{\partial\tau_{N-1}}(V(f(\tau_{N-1})))=\frac{\partial}{\partial\tau_{N-1}}(V(x_{N}))$
$=\frac{\partial V(x_{N})}{\partial x_{N}}\frac{\partial x_{N}}{\partial\tau_{N-1}}=\frac{\partial V(x_{N})}{\partial x_{N}}\frac{\partial f(\tau_{N-1})}{\partial\tau_{N-1}}$
$=\frac{\partial V(x_{N})}{\partial x_{N}}\begin{bmatrix}A_{N-1},B_{N-1}\end{bmatrix},$ (12)
where $\begin{bmatrix}A_{N-1},B_{N-1}\end{bmatrix}$ are the matrices associated to the linearized dynamics at timestep $N-1$, and $x_{N}=f(\tau_{N-1})$, $\frac{\partial V(x_{N})}{\partial x_{N}}$ is the gradient of the value function with respect to its inputs. For $H_{V}$:
$H_{V}$ $=\frac{\partial}{\partial\tau_{N-1}}(\Delta_{V})$
$=\frac{\partial}{\partial\tau_{N-1}}\left(\frac{\partial V(x_{N})}{\partial x_{N}}\right)\frac{\partial f(\tau_{N-1})}{\partial\tau_{N-1}}$
[…]
$=\frac{\partial x_{N}^{T}}{\partial\tau_{N-1}}\frac{\partial^{2}V(x_{N})}{\partial x_{N}^{2}}\frac{\partial x_{N}}{\partial\tau_{N-1}}$
$=\frac{\partial f(\tau_{N-1})^{T}}{\partial\tau_{N-1}}\frac{\partial^{2}V(x_{N})}{\partial x_{N}^{2}}\frac{\partial f(\tau_{N-1})}{\partial\tau_{N-1}}$
$=\begin{bmatrix}A_{N-1}\\ B_{N-1}\end{bmatrix}\frac{\partial^{2}V(x_{N})}{\partial x_{N}^{2}}\begin{bmatrix}A_{N-1},B_{N-1}\end{bmatrix},$ (13)
where we have assumed that the second derivative of the dynamics is zero, and $\frac{\partial^{2}V(x_{N})}{\partial x_{N}^{2}}$ is the Hessian of the value function with respect to its inputs.
In order to empirically show the relationship between the Hessian of the value function and what is learned in $Q_{N}$ and $p_{N}$, we perform a similarity study using linear probes. Particularly, we first train an AC-MPC agent to perform the stabilization at hover task, introduced in Section V-D. We choose this task here because, since the reward function is quadratic, it results in a optimization landscape that is easier to interpret. Since we need the derivative of the value function with respect to the full state, the critic needs to be trained with full state information. After training, we collect a dataset of uniformly distributed $5242882^{19}=$ 524288 datapoints, where every datapoint corresponds to a sample of $Q_{N}$, $p_{N}$, $H_{V}$ and $\Delta_{V}$, as per Eq. (V-E) and (V-E). Then, with this dataset, we train a single linear layer (wihtout non-linearity) to predict the elements of the learned $Q_{N}$ and $p_{N}$ matrices from the elements of the $H_{V}$ and $p_{V}$, respectively. In the hypothesized case where the information contained in the learned cost is related to the information encoded by the gradient and Hessian of the value function, we expect the regression to have a high accuracy.
In Fig. 9A, we show the true and predicted values of $Q_{N}$ and $p_{N}$. In order to show that the data fully correlates in all dimensions, the data is presented in a an lower triangular matrix for $Q$ and in a upper triangular matrix for $p$. Every row and column of this matrix is composed of the elements of the state space $x$, therefore showing the correlation from different 1-dimensional cuts – for the diagonal elements – and 2-dimensional cuts – for the rest. One can therefore see a high correlation between the predicted and the true values, therefore confirming a strong relationship between the value function and the learned cost parameters. An extra experiment was conducted where we have tried to infer the same Q and p values, but from a random dataset, as a validation that there is no overfitting behaviour happening. Indeed, in this case the linear classifier was not able to predict the true labels at all.
V-F How does our approach handle exploration?
Since our approach incorporates a model of the dynamics within the optimization problem, it has access to prior information before the training phase, and hence leverages an inductive bias in its architecture. This raises questions about the extent of exploration needed to discover an optimal policy and how the exploration parameter can be minimized to achieve optimal performance. To explore this, we study the effect of the initial policy standard deviation hyperparameter on the performance of both AC-MLP and AC-MPC. This hyperparameter influences the exploration behavior of the RL algorithms. In Fig. 10, we present reward plots for the training on the Vertical track under different values of this hyperparameter. The results show that AC-MLP requires a higher degree of exploration to identify the optimal solution. This indicates that AC-MPC effectively utilizes the prior dynamics knowledge, resulting in improved performance with reduced exploration. More interestingly, Fig. 10 shows that AC-MPC is less sensitive to the choice of this hyperparameter and yields high performance despite very low exploration values. On the other hand, AC-MLP fails to reach high-reward regions when trained with low values and generally underperforms in comparison to the proposed method.
Figure 10: Comparison of reward values during training for different exploration parameters [0.22, 0.3, 0.44, 0.6] for both AC-MLP and AC-MPC, for the Vertical track.
V-G How does our method perform in real-world environments?
We test our approach in the real world with a high-performance racing drone. In order to test simulation to real world transfer, we deploy the policy with two different race tracks: Circle track and SplitS track. We use the Agilicious control stack [124] for the deployment. The main physical parameters and components of this platform are based on [14], under the name 4s drone.
Figure 11: Top view of the performance of AC-MLP and AC-MPC policies deployed in both nominal and a realistic simulator for the SplitS track. The policies are tested for 64 different initial positions, distributed in a cube of $0.5m$ of side length.
@@ V-G How does our method perform in real-world environments?
Additionally, we perform a lap time comparison in the SplitS track. We train both AC-MPC and AC-MLP in the task of drone racing and execute different runs in two different simulation fidelities: nominal dynamics and realistic simulator. This realistic simulator, called NeuroBEM, and based in blade element momentum theory, comes from [122]. In Fig. 11 we show a top view of the deployment of both AC-MLP and AC-MPC in both of the aforementioned simulators. The policies are tested for 64 different initial positions, contained in a cube of $0.5$ m of side length. Additionally, Fig. 12 illustrates trajectories flown in two different tracks (Circle track and SplitS track) in simulation and the real world. Not only can our policy transfer to the real world without any fine-tuning, it can do so while at the limit of handling, being able to achieve speeds of up to $21$ $m/s$. Finally, in Table II, we show the results in terms of lap time and success rate of this comparison, including both simulation modalities, and the real-world results.
TABLE II: Performance comparison between AC-MLP and AC-MPC in the SplitS track
AC-MLP AC-MPC
Drone Model Lap Time [s] Success Rate [%] Lap Time [s] Success Rate [%]
Nominal 5.09 $\pm$ 0.008 100.0 5.13 $\pm$ 0.008 100.0
Realistic 5.179 $\pm$ 0.01 100.0 5.24 $\pm$ 0.01 100.0
Real World 5.39 $\pm$ 0.08 85.7 5.4 $\pm$ 0.082 87.5
As observed in the lap time results, both AC-MPC and AC-MLP exhibit similar performance in terms of lap time, indicating that they are largely on par in this aspect. This similarity is expected, given that their asymptotic performance and final reward values are quite close, as illustrated in Figure 4. However, it is important to note that our approach not only achieves on par performance but also demonstrates additional robustness to out-of-distribution scenarios (Figure 5) and changes in the system dynamics (Figure 6). This indicates that our method maintains its effectiveness under a wider range of conditions and uncertainties, while still matching the performance of the neural network-based architecture.
V-H How do training and inference times compare to standard actor-critic architectures?
In Table II, we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths.
TABLE III: Solve times and inference times for different variations.
@@ V-H How do training and inference times compare to standard actor-critic architectures?
VI Discussion and Conclusion
This work presented a new learning-based control framework that combines the advantage of differentiable model predictive control with actor-critic training. Furthermore, we showed that AC-MPC can leverage the prior knowledge embedded in the system dynamics to achieve better training performance (sample efficiency, robustness to hyperparameters) than the AC-MLP baseline, and to better cope with out-of-distribution scenarios, and with variations in the nominal dynamics, without any further re-training. Through empirical analysis, the interpretable nature of the differentiable MPC embedded in the actor-critic architecture allows us to conduct and empirical analysis that revealed a relationship between the learned value function and the learned cost functions. This provides a deeper understanding of the interplay between the reinforcement learning (RL) and MPC, offering valuable insights for future research. Additionally, our approach achieved zero-shot sim-to-real transfer, demonstrated by successfully controlling a quadrotor at velocities of up to 21 m/s in the physical world, being on par in performance with the AC-MLP baseline. In general, we show that our method can tackle challenging control tasks and achieves robust control performance for agile flight.
However, there are some limitations to be mentioned and to be improved in the future. First, an analytic model of the system is required for the differentiable MPC block, which limits our approach to mainly systems where the dynamics are known beforehand. Furthermore, training AC-MPC takes significantly longer than AC-MLP (see Table III), due to the fact that an optimization problem needs to be solved for both the forward and the backward pass through the actor network. In fact, there are open-source libraries [101, 102] that are recently evolving and implementing more efficient versions of differentiable MPC. Showcasing the approach in different tasks with different robots is another future direction.
We believe that our method represents an important step in the direction of generalizability and robustness in RL. The proposed method demonstrates that modular solutions, which combine the advantages of learning-centric and model-based approaches, are becoming increasingly promising. Our approach potentially paves the way for the development of more robust RL-based systems, contributing positively towards the broader goal of advancing AI for real-world robotics applications.
@@ Acknowledgments
We would like to thank Brandon Amos, for sharing his insights regarding the differentiable MPC code. We would also like to thank Jiaxu Xing, Leonard Bauersfeld and Alex Barden for the insightful discussions and support with the hardware setup.